• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name T. Chang

    Log entry time 08:00:49 on April19,2001

    Entry number 60783

    keyword=Owl shift summary

    The ep elastic scan night.

    00:06 start configuration change:
    Pr = -1.6395, Pl = 3.1563, L_angle=22.89, R_angle=48.5. Trigger downloaded. Waiting for the right arm dipole to ramp down.

    01:58 Finished configuration change.
    Run 3671 15cm dummy target, raster on, 30 uA
    Run 3672 9-foil carbon, raster off, 70uA.
    Run 3673 15cm H2, raster on, 70uA. After 5 min, the beam is off and MCC wants to do a QE measurement.
    Run 3674 15cm H2, 70uA

    03:05 Start configuration change.
    Pr = -1.5152, Pl = 3.2858, L_angle=21.31, R_angle=52.0, Trigger downloaded.
    Waiting for the right arm dipole to ramp down.
    03:31 Finish configuration changes.
    Run 3675 15cm dummy target, raster on, 30 uA.
    Run 3676 9-foil carbon, rastor off, 70 uA.
    Run 3677 15cm H2 target, raster on, 70uA.

    04:21 Start configuration changes
    Pr = -1.3593, Pl=3.4476, L_angle=19.28, R_angle=56.981, Trigger downloaded.
    Waiting for the right arm dipole.
    04:50 Finish configuration changes
    Run 3678 15cm dummy target, raster on, 30 uA.
    Run 3679 9-foil carbon, raster off, 70 uA.
    Run 3680 15cm H2 target, raster on, 70 uA.

    05:50 Start configuration changes. Bogies_LEFT died. push the green reset buttom.
    That fixed the problem.
    Pl=2.377, Pr=2.377, L_angle=32.762, R_angle=32.842, Trigger downloaded. waiting for the right arm dipole.
    07:09 Finished configuration changes
    Run 3683 15 cm dummy target, raster on, 30 uA.
    Run 3684 9-foil carbon, raster off, 70 uA.
    Run 3685 15cm H2 target, raster on, 70 uA.

    07:53 Finished one polarity measurement. Ramping down the magnets. Hand over to the next shift.